# Leveraging AI-Generated Emotional Self-Voice to Nudge People towards their Ideal Selves

### v1

@@ 5. Results and Analysis
We begin by describing the methods for analyzing the quantitative and qualitative results. We first present the results of our main hypotheses of the effect of the intervention on individuals’ affective states and perceptions towards their goals by comparing the results before and after the intervention. Then, we compare results between the conditions to dissect the effect of the individual components of the intervention.
@@ 5.1. Analysis methods
For the quantitative data, we used a linear mixed effects (LME) model (lme4 package in R (Bates et al., 2015)) to account for the repeated measurement nature of the data, namely each subject has 2 observations: pre-intervention and post-intervention. We used the random intercept model, allowing each subject to have a unique intercept. The predictors of interest are Test (pre vs post) and Condition (Control, Text, and ESV), and we control for demographic information like gender and age, as well as other aspects related to the intervention like vividness of mental imagery, Alexithymia, self-compassion, and self-discrepancy. The main outcomes are valence, arousal, resilience, confidence, motivation, and commitment. Intuitively, these outcomes should not be independent as outcomes from the same person have the same underlying determinants. However, we do not model the correlations among the outcomes at this point. Thus, we fit a random intercept linear mixed effect model for each outcome separately. Note that we encountered collinearity in the condition-related fixed effects for the outcomes confidence, motivation, and commitment. This was due to a near-perfect separation in the pre- and post-intervention data between the correlated conditions. To addrfor these outcomes, we instead conducted a two-way ANOVA on data that was Aligned Rank Transformed (Wobbrock et al., 2011), as the scores of the questions did not have a normal distribution (Shapiro-Wilk normality test p¡0.001). To identify the levels of our factors that were contributing to observed differences, we conducted a post-hoc pairwise comparison using the emmeans package in R (Searle et al., 1980), which produces an adjusted p-value.
@@
5.2. Comparison between pre- and post- intervention
5.2.1. Affective states in terms of valence and arousal
Overall, there is a significant increase (p¡.001) in valence ratings across most conditions for both scenarios, but the trends in the change in arousal ratings vary (Figure 6 top row). Participants overall felt unpleasant upon describing the negative scenarios related to goals pre-intervention and felt more positive after the intervention. Note that across conditions, Scenario 1 “failure of achieving a goal” has a lower pre-intervention valence score compared to Scenario 2 “difficulty maintaining a habit” (Scenario 1: $\mu$ = 2.91, $\sigma$ = 0.81; Scenario 2: $\mu$ = 4.16, $\sigma$ = 0.81). For Scenario 1, the LME model also revealed self-compassion and self-discrepancy had a significantly negative effect on arousal (self-compassion: estimate = -1.08, p¡ 0.001; self-discrepancy: estimate = -0.28, p¡ 0.01).
The VADER compound scores indicate that for both scenarios, once again, all three conditions show a transition from negative or neutral sentiment to a strong positive sentiment (Table 1). Scenario 2 (forming habits) generally elicited more positive responses across all conditions compared to Scenario 1 (past failures). The ESV condition appears to be the most effective in improving emotional states across both scenarios.
@@ 5.2.2. Resilience
For Scenario 1, we looked at participants’ resilience towards the failure of achieving a goal. Overall, participants across conditions had a significant (p¡0.05) increase in resilience after the intervention (Figure 6, bottom left). The LME model revealed that self-compassion had a significantly positive effect on resilience (estimate = 0.62, p¡ 0.001). This shows that the more self-compassion one has, the more resilient they are post-intervention.
We analyzed the themes of participants’ responses to the question “what was the reason you didn’t succeed?” While all three conditions predominantly attribute failure to internal factors, the ESV condition emphasizes personal shortcomings and emotional states: “A lack of encouragement, time, and especially self confidence” (P46), “I wasn’t in the right headspace” (P50), and “mindset and lack of commitment” (P60), while Text and Control conditions focus on external factors: “There were outside factors out of my control that contributed to regaining the weight” (P5), and “The reason I didn’t succeed was due to unexpected delays and ineffective time management” (P28). On the other hand, the ESV condition appeared to have more negative residual feelings and focus on current feelings whereas the Text and Control conditions focus more on future-oriented thinking.
The VADER analysis of the free responses shows a similar trend where the ESV condition has the least positive sentiment when describing their feeling about the failure post-intervention (Table 2 Scenario 1). On the other hand, it seems like the ESV condition had less negative sentiment when describing failure attribution post-intervention.
@@
5.2.3. Confidence and Motivation
Analyzing participants’ responses to the question “How do you feel now about your ability to establish the personal habit?”, the ESV group demonstrates more consistent positive feelings and self-belief: “I believe in myself” (P55), “It’s hard to do but I am going to keep trying’’ (P58), “I feel more capable of taking it on” (P41). On the other hand, the Text and Control groups expressed less certainty: “I think it can be achieved” (P25), “I feel like I need more exercise, but I feel so limited” (P5), and more measured optimism and more awareness of the challenge: “I feel mildly overwhelmed but cautiously optimistic that I can achieve this goal” (P40), “I feel optimistic but recognize the challenges” (P20). The VADER analysis of the free responses also shows the ESV condition has the most positive sentiment when describing their feeling about their capability of habit forming (Table 2 Scenario 2, Feeling).
Figure 7. Between-condition comparison of post-intervention ratings of Vividness, Similarity, and Positive Affect towards the Ideal Self .
@@ 5.2.4. Commitment
To summarize, the Text and ESV groups seem to have more commitment to the goal of establishing a personal habit, and they were more confident that the goal was achievable and needed positive reinforcement for motivation, whereas the Control group leaned towards pessimism with a focus on the challenges and not the opportunities.
@@ 5.3. Comparing Effectiveness of Ideal Self Representation Between Conditions
We measured the distance between the individual’s current self and their ideal self using the adapted version of the Future-Self Continuity Questionnaire, where we replaced “future-self” with “ideal-self”. Overall the post-intervention ratings of the three aspects: vividness, similarity, and preference (having a positive affect) are above neutral (score = 3.5) (Figure 7).
We did not observe any significant differences between the conditions in the quantitative results (Figure 7), but we found more nuanced differences in the qualitative responses. The Control condition group expressed a positive yet low intensity attitude towards the ideal self concept and felt a sense of distance: “I feel happy thinking about my ’ideal self’ as it’s something that everyone hopes to achieve. I don’t really feel that strongly about it since it seems distant to me” (P2), “The imagining did not invoke too much internal emotion” (P14). The Text condition group responded more strongly towards the generated ideal-self responses and appreciated the affirmation from a first-person perspective: “I really liked how the AI spoke in first person. It helped me feel like I was saying in affirmation to myself and that I could actually achieve what was written” (P26). Similarly, the ESV condition felt a strong sense of positivity and empathy from a first-person’s perspective: “It was encouraging to hear a generated response using my voice to hear my goals along with adjectives describing myself” (P56). “Its like hearing myself speak positively to me helps me be more positive” (P55). The generated ideal-self voice activated individuals: “…It was exactly what I needed to hear” (P59), “It made me enthusiastic enough to come up with a plan on how to undertake the problem and generate a working solution” (P53). More concretely, participants in the ESV condition explicitly mention being able to more vividly visualize an ideal version of themselves: “made me visualize a version of myself that is a bit more positive” (P48), “I have a positive opinion of the voice and text from the scenario-in fact, can even envision her in my mind (she doesn’t look like me, doesn’t have the same job as me…) but I like how articulate ‘I’ was and would feel good about myself saying what ‘she’ said” (P52).
@@ 6. System Characterizations
The goal of the characterization is to confirm our pipeline’s ability to generate the intended emotion. The output of our pipeline consists of text and speech output, and both contribute to the overall sentiment and attitude of the text. Complementary to the user study results which provide a subjective, qualitative evaluation of the emotional quality of the output, the technical validation tests are meant to provide a quantitative, objective rating of the emotional quality of the generation.
6.1. Validating the sentiment of the text content
We again chose VADER (Hutto and Gilbert, 2014) to quantitatively analyze the sentiment of the generated texts. Overall, the results show that generated ideal-self text responses are positive (Table 3), and more positive for Scenario 2 than Scenario 1, likely due to the nature of the context of the Scenario 2 focusing on future habits rather than past failures.
We compared the words participants used to describe the ideal self against their perceived attitude of the generated text. In both scenarios, the generated text seems to have been perceived as more overtly positive and emotionally supportive than the ideal self characteristics. The ideal self words tend to be more focused on personal qualities and actions, while the text descriptions often reflect the emotional impact of the generated content. In addition, the ideal self words are more action-oriented (e.g., “disciplined”, “consistent”, “organized”), while the text descriptions focus more on emotional states and motivation. Some key ideal self characteristics like “disciplined” and “hardworking” do not appear in the text descriptions, suggesting the generated text might not have emphasized these aspects as much. There is a greater alignment between ideal self and text descriptions in Scenario 2. In both cases, the generated text seems to have successfully captured many aspects of what people aspire to be, particularly in terms of positivity, resilience, and motivation.
@@ 6.1. Validating the sentiment of the text content
Table 4. Participants’ ratings of various aspects of the ESV speech.
@@ 6.2. Validating the audio quality of the speech
For the ESV condition, we asked participants questions about hearing their own voice and how they rate the generated self-voice. These participants had limited to no experience with voice cloning prior to the study ($\mu$ = 1.15, $\sigma$ = 0.49, 1=‘Once’, 2=‘A few times a year’). On average, they had a neutral to slightly negative feeling towards listening to their own voice ($\mu$ = 2.85, $\sigma$ = 1.31, 2=‘Slightly no’, 3=‘Indifferent’). Overall, these participants rated the generated self-voices to be familiar, similar to their own, natural, and have good quality (Table 4). In contrast to reactions to listening to their true voice, overall no participants responded negatively to the generated self-voice. Most reported being pleasantly surprised by the resemblance of the voice: “Caught off guard with how striking the resemblance is” (P54), “I couldn’t believe how quickly it cloned my voice and how accurate it was at reflecting my tone. Made me excited for the future” (P55), “Really got close to what I actually sound like” (P54).
Figure 8. emotion2vec results of the emotions represented in the generated ideal self speeches.
6.3. Validating the emotional quality of the speech
Once again, we compared the words participants used to describe the ideal self against their perceived attitude of the generated audio. In both scenarios, the generated voice seems to have been perceived as more emotionally nuanced and expressive compared to the more action-oriented ideal self characteristics. The voice appears to have successfully conveyed a sense of positivity and confidence that aligns well with people’s ideal self-perceptions. The realism and accuracy of the voice were frequently mentioned, suggesting that it felt authentic to listeners despite being generated. Some respondents found the voice to lack emotion or sound monotone, indicating that the emotional expressiveness might not have been consistent for all listeners. The voice descriptions include words like “calm” and “balanced”, which do not directly appear in the ideal self words but may reflect a desired state of being focused and organized. The voice was often described as “expressive”, which seems at times to be perceived as more overtly optimistic.
We chose the state-of-the-art emotion2vec model (Ma et al., 2023) to analyze the sentiment of the speech. We processed the generated speech for the ESV conditions and we analyzed at a sentence level. Figure 8 and Table 5 show the emotion2vec results for each sentence. Overall we see that the emotional representations of the two scenarios are distinct: with sentences in Scenario 2 sounding mostly “happy” while ones in Scenario 1 sounding “happy” but more “neutral”.
@@ 6.3. Validating the emotional quality of the speech
Table 5. emotion2vec results of the generated ESV speeches
@@ 6.4. Parameterization
Table 6 shows a summary of the final chosen parameters of the generated text and voice. Overall, the mean values are all above 0 (on a scale of -3 to +3), indicating that on average people prefer more positivity and expressiveness in the generated responses. However, these values have significant variance, pointing to the nuanced differences among individual preferences. We then contextualize the chosen parameter values by analyzing participants’ free responses to explaining their chosen parameters. Many participants chose parameters that aligned with their emotional state or how they wanted to feel about the situation. Some aimed for positivity to make themselves feel better, while others chose more negative sentiments to match their experience. Many participants in both conditions tried to find a balance between different aspects (e.g., positivity vs. realism, expressivity vs. calmness) to create the most satisfying output. For Scenario 2, participants were generally more focused on creating positive, motivational outputs, likely due to the forward-looking nature of forming a new habit. The parameters were often chosen with the specific goal of habit formation in mind, aiming for a tone that would be most conducive to achieving the goal.
@@ 7. Discussion
We first begin by discussing immediate findings based on the results of the study, and then we discuss the broader implications for theories related to ideal-self and goal-oriented behavior change and implications for designing future intervention systems.

### v2

@@ 5. Results and Analysis
We begin by describing the methods for analyzing the quantitative and qualitative results. We first present the results of our main hypotheses of the effect of the intervention on individuals’ affective states and perceptions towards their goals by comparing the results before and after the intervention. Then, we compare results between the conditions to dissect the effect of the individual components of the ESV intervention.
@@ 5.1. Analysis methods
For the quantitative data, we used a linear mixed effects (LME) model (lme4 package in R (Bates et al., 2015)) to account for the repeated measurement nature of the data, namely each subject has 2 observations: pre-intervention and post-intervention. We used the random intercept model, allowing each subject to have a unique intercept. The predictors of interest are Test (pre vs post) and Condition (Control, Text, and ESV), and we control for demographic information like gender and age, as well as other aspects related to the intervention like vividness of mental imagery, Alexithymia, self-compassion, and self-discrepancy. The main outcomes are valence, arousal, resilience, confidence, motivation, and commitment. Intuitively, these outcomes should not be independent as outcomes from the same person have the same underlying determinants. However, we do not model the correlations among the outcomes at this point. Thus, we fit a random intercept linear mixed effect model for each outcome separately. Note that we encountered collinearity in the condition-related fixed effects for the outcomes confidence, motivation, and commitment. This was due to a near-perfect separation in the pre- and post-intervention data between the correlated conditions. To address this, for these outcomes, we instead conducted a two-way ANOVA on data that was Aligned Rank Transformed (Wobbrock et al., 2011), as the scores of the questions did not have a normal distribution (Shapiro-Wilk normality test p¡0.001). To identify the levels of our factors that were contributing to observed differences, we conducted a post-hoc pairwise comparison using the emmeans package in R (Searle et al., 1980), which produces an adjusted p-value.
@@
Figure 7. Between-condition comparison of post-intervention ratings of Vividness, Similarity, and Positive Affect towards the Ideal Self. Error bars: SE.
5.2. Comparison between pre- and post- intervention
5.2.1. Affective states in terms of valence and arousal
Overall, the valence ratings became significantly (p¡.001) more positive across most conditions for both scenarios, but the trends in the change in arousal ratings vary (Figure 6 top row). Participants overall felt unpleasant upon describing the negative scenarios related to goals pre-intervention and felt more positive after the intervention. Note that across conditions, Scenario 1 “failure of achieving a goal” had a more negative pre-intervention valence score compared to Scenario 2 “difficulty maintaining a habit” (Scenario 1: $\mu$ = 2.91, $\sigma$ = 0.81; Scenario 2: $\mu$ = 4.16, $\sigma$ = 0.81). For Scenario 1, the LME model also revealed self-compassion and self-discrepancy had a significantly negative effect on arousal (self-compassion: $\beta$ = -1.08, p¡ 0.001; self-discrepancy: $\beta$ = -0.28, p¡ 0.01).
The VADER compound scores indicate that for both scenarios, once again, all three conditions show a transition from negative or neutral sentiment to a strong positive sentiment (Table 1). The ESV condition appears to be the most effective in improving emotional states across both scenarios.
@@ 5.2.2. Resilience
For Scenario 1, we looked at participants’ resilience towards failure of achieving a goal. Overall, participants across conditions had a significant (p¡0.05) increase in resilience after the intervention (Figure 6, bottom left). The LME model revealed that self-compassion had a significantly positive effect on resilience ($\beta$ = 0.62, p¡ 0.001). This shows that the more self-compassion one has, the more resilient one is post-intervention.
We analyzed the themes of participants’ post-intervention responses to the question “what was the reason you didn’t succeed?”. While all three conditions predominantly attribute failure to internal factors, the ESV condition emphasized personal shortcomings and emotional states: “A lack of encouragement, time, and especially self confidence” (P46), “I wasn’t in the right headspace” (P50), and “mindset and lack of commitment” (P60), while Text and Control conditions focused on external factors: “There were outside factors out of my control that contributed to regaining the weight” (P5), and “The reason I didn’t succeed was due to unexpected delays and ineffective time management” (P28). On the other hand, the ESV condition appeared to have more negative residual feelings and focused on current feelings whereas the Text and Control conditions focused more on future-oriented thinking.
The VADER analysis of the free responses shows a similar trend where the ESV condition had the least positive sentiment when describing their feeling about the failure post-intervention (Table 2 Scenario 1, Feeling). On the other hand, it seems like the ESV condition had less negative sentiment when describing failure attribution post-intervention (Table 2 Scenario 1, Reason).
5.2.3. Confidence and Motivation
Analyzing participants’ responses to the question “How do you feel now about your ability to establish the personal habit?”, the ESV group demonstrated more consistent positive feelings and self-belief: “I believe in myself” (P55), “It’s hard to do but I am going to keep trying’’ (P58), “I feel more capable of taking it on” (P41). On the other hand, the Text and Control groups expressed less certainty: “I think it can be achieved” (P25), “I feel like I need more exercise, but I feel so limited” (P5), and more measured optimism and more awareness of the challenge: “I feel mildly overwhelmed but cautiously optimistic that I can achieve this goal” (P40), “I feel optimistic but recognize the challenges” (P20). The VADER analysis of the free responses also shows the ESV condition had the most positive sentiment when describing their feeling about their capability of habit forming (Table 2 Scenario 2, Feeling).
@@ 5.2.4. Commitment
5.2.5. Clustering of Responders and Non-Responders
In addition to the main hypotheses and aforementioned results from the LME model, we tried to understand the difference between responders and non-responders to this type of cognitive intervention. We performed a hierarchical clustering of participants of all conditions. We found that for scenario 1, people who were the least respondent to the intervention were older and had higher self-compassion, whereas the most respondent group on average had a noticeably younger age, lower self-compassion, and a higher self-discrepancy. This makes sense intuitively, as people who are already self-compassionate might not need the intervention. For scenario 2, we found that people who had higher alexithymia did not respond to the intervention, which could be that they were less susceptible to the “self pep-talk”. The additional results are provided in the Appendix C.2.
@@ 5.3. Comparing Effectiveness of Ideal Self Representation Between Conditions
We measured the distance between the individual’s current self and their ideal self using the adapted version of the Future-Self Continuity Questionnaire, where we replaced “future-self” with “ideal-self”. Overall, the post-intervention ratings of the three aspects: vividness, similarity, and positive affect were above neutral (score = 3.5) (Figure 7).
We did not observe any significant differences between the conditions in the quantitative results (Figure 7), but we found more nuanced differences in the qualitative responses. The Control condition group expressed a positive yet low intensity attitude towards the ideal self concept and felt a sense of distance: “I feel happy thinking about my ’ideal self’ as it’s something that everyone hopes to achieve. I don’t really feel that strongly about it since it seems distant to me” (P2), “The imagining did not invoke too much internal emotion” (P14). The Text condition group responded more strongly towards the generated ideal-self responses and appreciated the affirmation from a first-person perspective: “I really liked how the AI spoke in first person. It helped me feel like I was saying in affirmation to myself and that I could actually achieve what was written” (P26). Similarly, the ESV condition felt a strong sense of positivity and empathy from a first-person’s perspective: “It was encouraging to hear a generated response using my voice to hear my goals along with adjectives describing myself” (P56). “Its like hearing myself speak positively to me helps me be more positive” (P55). The generated ideal-self voice activated individuals: “…It was exactly what I needed to hear” (P59), “It made me enthusiastic enough to come up with a plan on how to undertake the problem and generate a working solution” (P53). More concretely, participants in the ESV condition explicitly mentioned being able to more vividly visualize an ideal version of themselves: “made me visualize a version of myself that is a bit more positive” (P48), “I have a positive opinion of the voice and text from the scenario-in fact, can even envision her in my mind (she doesn’t look like me, doesn’t have the same job as me…) but I like how articulate ‘I’ was and would feel good about myself saying what ‘she’ said” (P52).
@@ 6. System Characterizations
The goal of the characterizations is to confirm our pipeline’s ability to generate the intended emotion. The output of our pipeline consists of text and speech output, and both contribute to the overall sentiment and attitude of the text. Complementary to the user study results which provide a subjective, qualitative evaluation of the emotional quality of the output, the technical validation tests are meant to provide a quantitative, objective rating of the emotional quality of the generation.
@@
6.1. Validating the sentiment of the text content
We again chose VADER (Hutto and Gilbert, 2014) to quantitatively analyze the sentiment of the generated texts. Overall, the results show that generated ideal-self text responses were positive (Table 3), and more positive for Scenario 2 than Scenario 1, likely due to the nature of the context of the Scenario 2 focusing on future habits rather than past failures.
We compared the words participants used to describe the ideal self against their perceived attitude of the generated text. In both scenarios, the generated text seems to have been perceived as more overtly positive and emotionally supportive than the participant-provided ideal self characteristics. The ideal self words tend to be more focused on personal qualities and actions, while the generated texts often reflect the emotional impact of the generated content. In addition, the ideal self words are more action-oriented (e.g., “disciplined”, “consistent”, “organized”), while the text descriptions focus more on emotional states and motivation. Some key ideal self characteristics like “disciplined” and “hardworking” do not appear in the text descriptions, suggesting the generated text might not have emphasized these aspects as much. There is a greater alignment between the ideal self and the generated text descriptions in Scenario 2. In both cases, the generated text seems to have successfully captured many aspects of what people aspire to be, particularly in terms of positivity, resilience, and motivation.
@@ 6.1. Validating the sentiment of the text content
Table 4. Participants’ ratings of various aspects of the ESV speech on a scale of 1 to 5 where a higher rating is better.
@@ 6.2. Validating the audio quality of the speech
For the ESV condition, we asked participants questions about hearing their own voice and how they would rate the generated self-voice. These participants had limited to no experience with voice cloning prior to the study ($\mu$ = 1.15, $\sigma$ = 0.49, 1=‘Once’, 2=‘A few times a year’). On average, they had a neutral to slightly negative feeling towards listening to their own voice ($\mu$ = 2.85, $\sigma$ = 1.31, 2=‘Slightly no’, 3=‘Indifferent’). Overall, these participants rated the generated self-voices to be familiar, similar to their own, natural, and have good quality (Table 4). In contrast to reactions to listening to their true voice, overall no participants responded negatively to the generated self-voice. Most reported being pleasantly surprised by the resemblance of the voice: “Caught off guard with how striking the resemblance is” (P54), “I couldn’t believe how quickly it cloned my voice and how accurate it was at reflecting my tone. Made me excited for the future” (P55), “Really got close to what I actually sound like” (P54).
@@
Table 5. Summary results of emotion classification (emotion2vec) of the generated ESV speeches
Figure 8. Emotion classification results (emotion2vec) of the emotions represented in the generated ESV speeches.
@@
6.3. Validating the emotional quality of the speech
Once again, we compared the words participants used to describe their ideal self characteristics against their perceived attitudes of the generated audio. In both scenarios, the generated voice seemed to have been perceived as more emotionally nuanced and expressive compared to the more action-oriented ideal self characteristics. The voice appeared to have successfully conveyed a sense of positivity and confidence that aligns well with people’s ideal self-perceptions. The realism and accuracy of the voice were frequently mentioned, suggesting that it felt authentic to listeners despite being generated. Some respondents found the voice to lack emotion or sound monotone, indicating that the emotional expressiveness might not have been consistent for all listeners. The voice descriptions included words like “calm” and “balanced”, which did not directly appear in the ideal self words but may reflect a desired state of being focused and organized. The voice was often described as “expressive”, which seemed at times to be perceived as overtly optimistic.
We chose the state-of-the-art emotion2vec model (Ma et al., 2023) to analyze the sentiment of the speech. We processed the generated speeches for the ESV conditions and we analyzed them at a sentence level. Figure 8 and Table 5 show the emotion2vec results. Overall we see that the emotional representations of the two scenarios are distinct: with sentences in Scenario 2 sounding mostly “happy” while ones in Scenario 1 sounding “happy” but more “neutral”.
@@ 6.4. Parameterization
Table 6 shows a summary of the participant-tuned parameters of the generated text and voice. Overall, the mean values are all above 0 (on a scale of -3 to +3), indicating that on average people prefer more positivity and expressiveness in the generated responses. However, these values have significant variance, pointing to the nuanced differences among individual preferences. We then contextualized the chosen parameter values by analyzing participants’ free responses to explaining their chosen parameters. Many participants chose parameters that aligned with their emotional state or how they wanted to feel about the situation. Some aimed for positivity to make themselves feel better, while others chose more negative sentiments to match their experience. Many participants in both conditions tried to find a balance between different aspects (e.g., positivity vs. realism, expressivity vs. calmness). For Scenario 2, participants were generally more focused on creating positive, motivational outputs, likely due to the forward-looking nature of forming a new habit. The parameters were often chosen with the specific goal of habit formation in mind, aiming for a tone that would be most conducive to achieving the goal.
@@ 7. Discussion
We begin by discussing immediate findings based on the results of the study, and then we discuss the broader implications for theories related to ideal-self and goal-oriented behavior change and implications for designing future intervention systems.